Guidelines for authors and submission template

Alan N. Other^{1}†, H. - C. Smith^{1, 2} and J. Q. Public^{2}

1Department of Chemical Engineering, University of America, Somewhere, IN 12345, USA 2Department of Aerospace and Mechanical Engineering, University of Camford, Academic

Street, Camford CF3 5QL, UK

(Received xx; revised xx; accepted xx)

This file contains instructions for authors planning to submit a paper to the Journal of Fluid Mechanics. These instructions were generated in L^{A}TEX using the JFM style, so the L^{A}TEX source file can be used as a template for submissions. The present paragraph appears in the abstract environment. All papers should feature a single-paragraph abstract of no more than 250 words, which provides a summary of the main aims and results.

1. How to submit to the Journal of Fluid Mechanics

Authors must submit using the online submission and peer review system, Scholar One Manuscripts (formerly Manuscript Central) at http://mc.manuscriptcentral.com/jfm. If visiting the site for the first time, users must create a new account by clicking on ‘register here’. Once logged in, authors should click on the ‘Corresponding Author Centre’, from which point a new manuscript can be submitted, with step-by-step instructions provided. Authors must at this stage specify whether the submission is a standard paper, or a JFM Rapids paper (see §4 for more details). In addition, authors must specify an editor to whom the paper will be submitted from the drop-down list provided. Note that all editors exclusively deal with either standard papers or JFM Rapids (clearly indicated on the list), so please ensure that you choose an editor accordingly. Once your submission is completed you will receive an email confirmation. Book reviews should not be submitted via the online submission site and should instead be submitted by email to c.p.caulfield@damtp.cam.ac.uk.

2. Rules of submission

Submission of a paper implies a declaration by the author that the work has not previously been published, that it is not being considered for publication elsewhere and that it has not already been considered by a different editor of the Journal. Note that a report on a conference must be submitted within 3 months of the meeting.

3. Types of paper

3.1. Standard papers

Regular submissions to JFM are termed ‘standard papers’. Note that such papers must contain original research, as review articles are not published in JFM. Papers should be written in a concise manner; though JFM has no page limit, each paper will be judged

† Email address for correspondence: jfm@damtp.cam.ac.uk

on its own merits, and those deemed excessive in length will be rejected or will require significant revision.

3.2. JFM Rapids

JFM Rapids is an evolution of the Fast Track scheme, and is devoted to the rapid publication of short, high-impact papers across the full range of fluid mechanics. Rapids must occupy 10 or fewer printed pages, and must be submitted in L^{A}TEX using the jfm.cls class file, so as to ensure that they meet the page limit and to expedite their production. The principal and over-riding objective is to publish papers of the highest scientific quality. Once a paper is submitted, referees are asked to provide reports with a short turnaround. In order to be accepted for publication in JFM Rapids, such papers must require only minor revisions to improve clarity, usually without recourse to a second round of refereeing. Where the editor, guided by the referees, judges that a paper has merit but requires substantial revision that will require significant refereeing, a decision of ‘revise and resubmit’ will be given. On re-submission, such papers will be handled as standard JFM papers and will not subsequently appear as JFM Rapids.

Rapids will be published online within one month of final acceptance. They will appear within a new devoted section on the Journal of Fluid Mechanics website, which has been re-designed as part of this development. Each Rapid will still be cited and indexed as a JFM article but with a distinctive Rapids identifier, and will be assigned to a JFM volume. Rapids will not be included in the regular print versions of JFM, but a collection of all Rapids will be available to order in print at the end of each year. JFM Rapids will have free colour and for the first year of operation will be freely accessible to all.

4. File types

Authors are strongly encouraged to compose their papers in L^{A}TEX, using the jfm.cls style file and supporting files provided at http://journals.cambridge.org/images/ fileUpload/documents/jfm-ifc.zip, with the jfm-instructions.tex file serving as a template (note that this is mandatory for JFM Rapids). A PDF of the L^{A}TEX file should then be generated and submitted via the submission site. Please note that PDFs larger than 3MB are not acceptable for review. There is no need to submit the L^{A}TEX source file alongside the PDF, but upon provisional acceptance of the paper, the L^{A}TEX source file, along with individual figure files and a PDF of the final version, will need to be submitted for typesetting purposes. Authors may also compose standard papers in Word, though this will lead to the paper spending a longer period in production. If using Word, please note that equations must NOT be converted to picture format and the file must be saved with the option ‘make equation editable’.

5. Preparing your manuscript

Authors should write their papers clearly and concisely in English, adhering to JFM’s established style for notation, as provided in §??. We encourage the submission of online supplementary material alongside the manuscript where appropriate (see section ??). Metric units should be used throughout and all abbreviations must be defined at first use, even those deemed to be well known to the readership. British spelling must be used, and should follow the Shorter Oxford English Dictionary.

Figure 1.

5.1. Figures

Figures should be as small as possible while displaying clearly all the information required, and with all lettering readable. Every effort should be taken to avoid figures that run over more than one page. Figures submitted in colour will appear online in colour but, with the exception of JFM Rapids, all figures will be printed in black and white unless authors specify during submission that figures should be printed in colour, for which there is a charge of £300 for the first page and £150 for each subsequent page that uses colour (colour is free for JFM Rapids). Note that separate figures for online and print will not be accepted and it is the author’s responsibility to ensure that if a figure is to appear in colour online only, that same figure will still be meaningful when printed in black and white (for example, do not rely upon colours to distinguish lines if those colours will just appear as similar shades of grey when printed). If using colour, authors should also ensure that some consistency is applied within the manuscript. For review purposes figures should be embedded within the manuscript. Upon final acceptance, however, individual figure files will be required for production. These should be submitted in EPS or high-resolution TIFF format (1200 dpi for lines, 300 dpi for halftone and colour in CMYK format, and 600 dpi for a mixture of lines and halftone). The minimum acceptable width of any line is 0.5pt. Each figure should be accompanied by a single caption, to appear beneath, and must be cited in the text. Figures should appear in the order in which they are first mentioned in the text and figure files must be named accordingly (‘Figure 1.eps’, ‘Figure 2a.tiff’, etc) to assist the production process (and numbering of figures should continue through any appendices). The word figure is only capitalized at the start of a sentence. For example see figures ?? and ??. Failure to follow figure guidelines may result in a request for resupply and a subsequent delay in the production process. Note that all figures are relabelled by the typesetter, so please ensure all figure labels are carefully checked against your originals when you receive your proofs.

4 A. N. Other et al

Table 1.